# The Mind in the Machine: A Survey of Incorporating Psychological Theories in LLMs

### v1

@@ Abstract
Psychological insights have long shaped pivotal NLP breakthroughs, including the cognitive underpinnings of attention mechanisms, formative reinforcement learning, and Theory of Mind-inspired social modeling. As Large Language Models (LLMs) continue to grow in scale and complexity, there is a rising consensus that psychology is essential for capturing human-like cognition, behavior, and interaction. This paper reviews how psychological theories can inform and enhance stages of LLM development, including data, pre-training, post-training, and evaluation&application. Our survey integrates insights from cognitive, developmental, behavioral, social, personality psychology, and psycholinguistics. Our analysis highlights current trends and gaps in how psychological theories are applied. By examining both cross-domain connections and points of tension, we aim to bridge disciplinary divides and promote more thoughtful integration of psychology into future NLP research.
@@ 1 Introduction
As Large Language Models (LLMs) grow in scale and complexity, the Natural Language Processing (NLP) community is increasingly recognizing psychology as essential for capturing human-like cognition, behavior, and interaction (Qu et al., 2024; Lewis, 2025). Psychology, grounded in empirically validated and computationally adaptable frameworks (Sartori & Orrù, 2023; Ong, 2024), can address core LLM challenges such as reasoning fidelity, context retention, and user interaction. Reflecting these strengths, psychological insights have long shaped pivotal NLP advances, including the cognitive inspirations of attention mechanisms, formative reinforcement learning approaches, and Theory of Mind-inspired social modeling.
@@
Despite extensive multidisciplinary efforts, a holistic review systematically integrating psychology across the LLM lifecycle remains missing. Most surveys and position papers remain fragmented, typically falling into three broad categories: (1) Some investigate how LLMs can empower traditional psychology or cognitive science research, for instance by modeling human reasoning and behavior at scale (Demszky et al., 2023; Abdurahman et al., 2024; Ong, 2024; Ke et al., 2024). (2) Others approach LLMs as subjects of psychological analysis, aiming to adapt or extend psychological theory – such as personality or cognition frameworks – to interpret and evaluate model behavior (Li et al., 2024b; Hagendorff et al., 2023; tse Huang et al., 2024; Pellert et al., 2024). (3) Finally, a third group leverages a single or limited set of psychological constructs to enhance model alignment or multi-agent frameworks – improving system reliability, social interaction, and trustworthiness (Liu et al., 2023; Dong et al., 2024b). This includes research on social influence for AI safety (Zeng et al., 2024a), moral reasoning in legal tasks (Almeida et al., 2024), and partial integrations of social or developmental psychology (Sartori & Orrù, 2023; Zhang et al., 2024c; Serapio-García et al., 2025). However, no existing work provides a unified map of how diverse psychological sub-areas can be harnessed, from data through application. Our survey fills this gap by offering a stage-wise view of how psychology can strengthen LLM capabilities and alignment across the entire lifecycle.
@@
To address this gap, we present a structured review that situates psychological theories from six major areas across the entire LLM development pipeline. Drawing on research from multiple NLP and AI conferences [^1] We survey 175 papers from major *CL venues (ACL Anthology), plus COLING, NeurIPS, ICML, ICLR, and influential arXiv preprints from late 2021 to early 2025 via keyword search., our contributions are twofold: (1) We systematically review psychological theories applied in key stages of LLM development, identifying gaps and inconsistencies. (2) We highlight under-explored concepts alongside critical issues and debates at the intersection of psychology and NLP. Collectively, these contributions demonstrate how integrating diverse psychological frameworks can strengthen LLM design, enhance alignment, and broaden the practical and ethical impact of modern NLP systems.
@@
As shown in Figure 1, the remainder of this paper illustrates how cognitive, developmental, behavioral, social, psycholinguistic, and personality theories integrate into four key stages of LLM development: preprocessing (Section 2), pre-training (Section 3), post-training (Section 4), and evaluation and application (Section 5). Finally, Section 6 examines three overarching questions: How does current LLM development leverage psychological theories? Which untapped psychological insights could advance LLM development? And what debates loom at the intersection of NLP and psychology?
@@
2 Preprocessing
We begin our stage-by-stage analysis of LLM development with preprocessing, the foundation that shapes downstream capabilities. During data collection and processing, psychology provides valuable frameworks into human learning and filtering, underscoring the need for realistic, developmentally informed datasets and effective filtering strategies.
Data Collection  Recent evidence shows that LLMs can align with human brain responses under biologically plausible training conditions (Hosseini et al., 2024), despite LLMs typically requiring orders of magnitude more training data than humans receive. This supports the application of ecological validity (Schmuckler, 2001) that emphasizes real-world data to mimic cognitive development. Drawing from this principle, Jagadish et al. (2024) selects linguistically diverse environments to reflect children’s language acquisition processes, while Feng et al. (2024) utilizes child-directed speech to mirror naturalistic learning contexts. In parallel, developmental theories like incremental numerical understanding (Piaget, 2013) that views numerical concepts as gradually acquired through exposure offer insights into sequential data organization. Applying this framework, Sharma et al. (2024) introduces mathematically coherent numeric anchors to align the collection process.
Data Preprocessing  Data preprocessing inspired by cognitive psychology involves refining data to enhance informational coherence prior to training. Selective attention (Treisman, 1969) – distinct from the attention mechanisms in transformers – prioritizes cognitively salient information while filtering out irrelevant stimuli. Implementing this principle, Nottingham et al. (2024) develops a preprocessing model that identifies and filters irrelevant data. Meanwhile, predictive coding (Rao & Ballard, 1999) proposes anticipatory processing based on prior knowledge, providing a framework for structuring sequential information. Leveraging this insight, Araujo et al. (2021) structures sentences to enable anticipation of subsequent content, improving semantic coherence through expectation-driven processing.
@@ 3 Pre-Training
Building on the foundations established during preprocessing, pre-training in LLMs mirrors aspects of human cognitive development, where core linguistic and reasoning abilities emerge through exposure to stimuli. This section explores how psychological insights inform observational learning and knowledge acquisition in LLMs.
Observational Learning and Self-Supervision
Incremental cognitive development (Piaget, 1976), which posits children acquire knowledge through sequential tasks, informs how LLMs can master nuanced concepts with explicit structured exposure. This principle manifests in Schulze Buschoff et al. (2023)’s gradually expanding pre-training tasks and Chen et al. (2024d)’s use of contradictory historical tasks for conceptual restructuring. Additionally, scaffolding theory (Park & Reuter-Lorenz, 2009), which emphasizes gradually challenging interactions, addresses the challenge of maintaining coherent learning trajectories – demonstrated by Borges et al. (2024)’s structured feedback loops and Sonkar et al. (2023)’s dynamic adjustment of task complexity based on performance metrics.
World Knowledge Acquisition
Semantic coherence during pre-training draw insights from top-down and bottom-up perception (Gregory, 1997), which frames cognition as interaction between conceptual frameworks and detailed data. Top-down processing is leveraged to prioritize high-level semantic processing before syntactic details (Rawte et al., 2022) and to generate test case (Zhang et al., 2024b). Meanwhile, Pang et al. (2023) fuses granular bottom-up encoding with top-down corrections for document summarization.
@@ 4 Post-Training and Alignment
With foundational knowledge acquired in pre-training, post-training refines LLMs from general proficiency to task-specific, goal-oriented behavior. We explore how psychology guide post-training setups for more context-aware, interpretable, and human-aligned outcomes.
Supervised Fine-Tuning (SFT)  In SFT, works that draw on psychological insights focus on retaining contextual information. Building on working memory theory (Baddeley & Hitch, 1974b) that proposes a short-term system for temporarily holding information, Kang et al. (2024) adds a module into Decision Transformers, enabling LLMs to retain and process short-term information. Beyond retaining, Li et al. (2023) proposes a working memory approach that dynamically balances stored information with provided contexts for SFT.
Reinforcement Learning from Human Feedback (RLHF) A classic behavioral theory, the Operant Conditioning theory posits that behaviors are systematically strengthened or weakened by the consequences (rewards or punishments) that immediately follow them (Thorndike, 1898; Skinner, 1957). The principles of reinforcement learning align closely with this psychological framework, particularly in the post-training phase of LLM development, where RLHF explicitly operationalizes Operant Conditioning theory to align model behaviors with human values and preferences. Through repeated feedback, the model gradually adapts to favor outputs that consistently yield higher reward signals, thereby aligning closer with human preferences – reflecting Thorndike’s Law of Effect, which asserts that behaviors followed by satisfying outcomes are more likely to recur (Lambert et al., 2023). During RLHF, the model generates responses, and a learned reward function $R(x)$ assigns scores to outputs $x$, guiding subsequent policy updates. For instance, Ouyang et al. (2022) train InstructGPT using Proximal Policy Optimization (Schulman et al., 2017), rewarding responses preferred by humans and penalizing less desirable ones. Foundational frameworks (Christiano et al., 2017; Sutton & Barto, 2018; Stiennon et al., 2022) established methods for explicitly translating human judgments into reward signals, operationalizing the insights of Operant Conditioning. More recent work incorporates human cognitive biases (Siththaranjan et al., 2024) and personalizes reward functions for individual values (Poddar et al., 2024). These developments illustrate how Operant Conditioning remains central to aligning LLMs with nuanced human values. While our survey focuses on psychological dimensions, a technical overview of RLHF methods is provided in Appendix B.
@@ 5 Evaluation and Application
In the evaluation and application stage of LLM development, psychology offers valuable tools for both assessing and enhancing model behavior. In this section, we review three key areas of challenges where psychology can inform this stage: (1) evaluating emergent capabilities such as reasoning, (2) improving task performance in domains involving human cognition, and (3) designing socially aware, multi-agent systems.
@@ 5.1.1 Social Reasoning and Intelligence
Notably, Theory of Mind (ToM) offers a framework for evaluating how individuals understand and attribute mental states – such as beliefs, desires, and intentions – to others (Baron-Cohen et al., 1985). By measuring an LLM’s capacity to represent and reason about beliefs, researchers can assess core social intelligence. Recent benchmarks include ToMBENCH (Chen et al., 2024c), OpenToM (Xu et al., 2024a), HI-TOM (Wu et al., 2023), and FANTOM (Kim et al., 2023), each probing distinct facets of ToM. Extending these efforts to spoken dialogues, Soubki et al. (2024) reveal lingering gaps between LLM and human performance. Surveys (Ma et al., 2023; Sap et al., 2022) then consolidate methods and underscore the challenges of building robust ToM-based evaluations.
Beyond individual cognition, social influence theories, such as Conformity Theories, capture how group pressure shapes individual judgments (Asch, 2016). Recent work tests LLM-based agents’ collaboration and bias dynamics under these principles (Zhang et al., 2024c; Jin et al., 2024b), bridging individual and group-level cognition.
Emotion recognition is another pillar of social intelligence. Ekman’s Basic Emotion Theory (Ekman, 1992) identifies six universal emotions, often used for labeling (Gong et al., 2024), while Dimensional Models like the Circumplex Model conceptualize emotions along valence and arousal (Morrill et al., 2024). LLMs increasingly excel at emotion recognition, benefitting dialogue and sentiment tasks (Zhang et al., 2024e; Wu et al., 2024c; d; Sabour et al., 2024).
@@ 5.1.2 Language Proficiency
Recent work adopts psycholinguistic assessments, originally designed for humans, to test LLMs’ language proficiency. These experiments probe a wide range of linguistic domains: morphology (Anh et al., 2024), syntax (Liu et al., 2024b; Hale & Stanojević, 2024), phonology (Duan et al., 2025), semantics (Duan et al., 2025; Hayashi, 2025) and their interactions (Miaschi et al., 2024; Zhou et al., 2025).
Although LLMs exhibit comparable performance to human speakers on many psycholinguistic tasks, the underlying processing mechanism they rely on may seem different from humans (Lee et al., 2024). Human language acquisition is often characterized by the Poverty of the Stimulus, i.e. children acquire complex grammar from relatively little input (Chomsky, 1980), whereas LLMs typically require developmentally implausible amounts of linguistic data to learn morphological rules. On the other hand, some evidence suggests that the learning patterns of LLMs mirror aspects of human language acquisition (Liu et al., 2024b).
Several studies have explored the pragmatic abilities of LLMs, motivated by the close link between language and broader cognitive functions in humans. Grice (1975)’s Theory of conversational implicature posits that utterance interpretation depends on both literal content and surrounding context. Researchers (Bender & Koller, 2020; Gubelmann, 2024) have contrasting perspectives on LLMs with respect to the Harnad (1990)’s Symbol Grounding Problem, i.e. linguistic symbols must be grounded in sensorimotor interactions to be meaningful. Failures of LLMs in pragmatic and semantic tasks (Kibria et al., 2024; Zeng et al., 2025), as well as their neuron patterns (Wu et al., 2024b), point to limitations beyond pure linguistic knowledge, which potentially parallel human higher-level cognitive processes.

### v2

# A Review of Incorporating Psychological Theories in LLMs

@@ Abstract
Psychological insights have long shaped pivotal NLP breakthroughs, from attention mechanisms to reinforcement learning and social modeling. As Large Language Models (LLMs) develop, there is a rising consensus that psychology is essential for capturing human-like cognition, behavior, and interaction. This paper reviews how psychological theories can inform and enhance stages of LLM development. Our review integrates insights from six subfields of psychology, including cognitive, developmental, behavioral, social, personality psychology, and psycholinguistics. With stage-wise analysis, we highlight current trends and gaps in how psychological theories are applied. By examining both cross-domain connections and points of tension, we aim to bridge disciplinary divides and promote more thoughtful integration of psychology into NLP research.
A Review of Incorporating Psychological Theories in LLMs
Zizhou Liu∗♠,     Ziwei Gong∗♠,     Lin Ai∗♠,     Zheng Hui∗†,     Run Chen∗♠, Colin Wayne Leach,     Michelle R. Greene,     Julia Hirschberg Equal contributions. Department of Computer Science, Columbia University, Department of Psychology, Barnard College, The Language Technology Lab, University of Cambridge {sara.ziweigong, lin.ai, runchen,julia}@cs.columbia.edu {cleach, mgreene}@barnard.edu,{zl2889, zh2483}@columbia.edu
@@ 1 Introduction
As Large Language Models (LLMs) grow in scale and complexity, the Natural Language Processing (NLP) community increasingly sees psychology as key to capturing human-like cognition, behavior, and interaction (Qu et al., 2024; Lewis, 2025). Psychology, grounded in empirically validated and computationally adaptable frameworks (Sartori and Orrù, 2023; Ong, 2024), offer principled tools for addressing persistent LLM challenges such as reasoning fidelity, context retention, and user interaction. Early intersections of psychology and NLP predate modern LLMs. Classic systems such as ELIZA Weizenbaum (1966) framed conversations around simple pattern matching, while early cognitive and psycholinguistic work influenced representation and processing assumptions. Today’s LLMs revive these threads at a new scale, but psychology is best viewed as one productive lens among others, complementary to linguistics, HCI, and sociology, rather than an inevitable or singular foundation. Reflecting these strengths, psychological insights have driven NLP advances, including the cognitive inspirations of attention mechanisms, formative reinforcement learning approaches, and social modeling for agents.
@@
Despite extensive multidisciplinary efforts, a holistic review systematically integrating psychology across the LLM lifecycle remains missing. Most surveys and position papers remain fragmented, typically falling into three broad categories: (1) Some investigate how LLMs can empower traditional psychology or cognitive science research, for instance by modeling human reasoning and behavior at scale (Abdurahman et al., 2024a; Demszky et al., 2023; Ong, 2024; Ke et al., 2024). (2) Others approach LLMs as subjects of psychological analysis, aiming to adapt or extend psychological theory, such as personality or cognition frameworks, to interpret and evaluate model behavior (Brito et al., 2025; Li et al., 2024b; Hagendorff et al., 2023; tse Huang et al., 2024; Pellert et al., 2024). (3) Finally, a third group leverages a single or limited set of psychological constructs to enhance model alignment or multi-agent frameworks – improving system reliability, social interaction, and trustworthiness (Dong et al., 2024b; Liu et al., 2023). This includes research on social influence for AI safety (Zeng et al., 2024a), moral reasoning in legal tasks (Almeida et al., 2024), and partial integrations of social or developmental psychology (Sartori and Orrù, 2023; Serapio-García et al., 2025; Zhang et al., 2024c). However, no existing work provides a unified map of how diverse psychological sub-areas can be harnessed, from data through application. Our survey fills this gap by offering a stage-wise view of how psychology can strengthen LLM capabilities and alignment across the entire lifecycle.
@@
To address this gap, we present a structured review that situates psychological theories from six major areas across the entire LLM development pipeline. The contributions of our survey [^1] We survey 227 papers from major *CL venues, plus COLING, NeurIPS, ICML, ICLR, and influential arXiv preprints. Appendix LABEL:appendix:keywords details paper selection process. are twofold: (1) We systematically review psychological theories applied in key stages of LLM development, identifying gaps and inconsistencies. (2) We highlight under-explored concepts alongside critical issues and debates at the intersection of psychology and NLP. Collectively, these contributions demonstrate how integrating diverse psychological frameworks can strengthen LLM design, enhance alignment, and broaden the practical and ethical impact of modern NLP systems.
@@
As shown in Figure 1, the remainder of this paper illustrates how cognitive, developmental, behavioral, social, psycholinguistic, and personality theories integrate into four key stages of LLM development: preprocessing (Section 1), pre-training (Section 3), post-training (Section 4), and evaluation and application (Section 5). Finally, Section 6 discusses three central questions: (i) How does current LLM development leverage psychological theories? (ii) Which untapped psychological insights could advance LLM development? (iii) What debates loom at the intersection of NLP and psychology?
2 Preprocessing
@@
We begin the stage-by-stage analysis of LLM development with preprocessing, the foundation that shapes downstream capabilities. Psychology provides valuable frameworks for understanding how humans acquire and filter information, underscoring the need for realistic, developmentally informed datasets and effective filtering strategies in data collection and processing.
Data Construction  Recent evidence shows that LLMs can align with human brain responses under biologically plausible training conditions (Hosseini et al., 2024), despite LLMs typically requiring orders of magnitude more training data than human. This supports the application of ecological validity (Schmuckler, 2001) that emphasizes real-world data to mimic cognitive development. To reflect children’s language acquisition processes, Jagadish et al. (2024) selects linguistically diverse environments, Feng et al. (2024) utilizes child-directed speech, while Nikolaus et al. (2022b) collects child cartoon. In parallel, incremental numerical understanding (Piaget, 1953) that views numerical concepts as gradually acquired through exposure is applied to sequential data collection with mathematically coherent numeric anchors (Sharma et al., 2024). Lastly, Reuben et al. (2025) provides a systematic framework to reformulate psychological questionnaires for LLM assessment.
Data Preprocessing  Data preprocessing inspired by cognitive psychology involves refining data to enhance informational coherence prior to training. Selective attention (Treisman, 1969), prioritizing salient information while filtering out irrelevant stimuli, was implemented to develop a preprocessing model that filters irrelevant data (Nottingham et al., 2024). Meanwhile, predictive coding proposing anticipatory processing based on prior knowledge (Rao and Ballard, 1999), was leveraged by Araujo et al. (2021) to enable anticipation of subsequent content, improving semantic coherence through expectation-driven processing. Lastly, drawing insights from knowledge acquisition of children, Ficarra et al. (2025) redefines lexical knowledge in data to capture distributional information based on target word.
@@ 3 Pre-Training
Building on the foundations established during preprocessing, pre-training mirrors human cognitive development, where linguistic and reasoning abilities emerge through exposure to stimuli. This section explores how psychology inform observational learning and knowledge acquisition in LLMs.
Observational Learning Incremental cognitive development (Piaget, 1976), which posits children acquire knowledge through sequential tasks, informs how LLMs can master nuanced concepts with explicit structured exposure. It is used as a motivating analogy for curriculum and sequence effects (Bengio et al., 2009) (i.e., how staged exposure can scaffold capability growth) rather than to claim a theoretical identity with next-token prediction. This principle manifests in Schulze Buschoff et al. (2023)’s gradually expanding pre-training tasks, Chen et al. (2024d)’s contradictory historical tasks and Ma et al. (2025c)’s trial-and-demonstration framework. In contrast, self-supervised objectives align more closely with use-based views of meaning, resonating with Wittgenstein’s claim that meaning emerges through use (Strawson, 1954). Additionally, scaffolding theory (Park and Reuter-Lorenz, 2009), which emphasizes gradually challenging interactions, informs maintaining coherent learning trajectories through Borges et al. (2024)’s structured feedback loops and Sonkar et al. (2023)’s dynamic task complexity.
Knowledge Acquisition Semantic coherence during pre-training draw insights from top-down and bottom-up perception (Gregory, 1997), which frames cognition as interaction between conceptual frameworks and detailed data. Top-down processing is leveraged to prioritize semantic processing before syntactic details (Rawte et al., 2022) and to generate test cases (Zhang et al., 2024b). Meanwhile, to enhance perception modeling, Pang et al. (2023) fuses bottom-up encoding with top-down corrections, and Nikolaus and Fourtassi (2021) models production-based learning. Introducing working memory theory (Baddeley and Hitch, 1974) that proposes a short-term system for temporarily holding information, Mita et al. (2025) simulates critical period dynamics with growing memory capacity to enhance performance.
@@ 4 Post-Training and Alignment
With foundational knowledge acquired in pre-training, post-training refines LLMs from general proficiency to task-specific behavior. We explore how psychology guide post-training for context-aware, interpretable, and human-aligned outcomes.
Supervised Fine-Tuning (SFT)  In SFT, works that draw on psychological insights focus on retaining and learn contextual information. Building on working memory theory, Kang et al. (2024) adds a working memory module to retain short-term information, while Li et al. (2023) dynamically balances memory with contexts to improve robustness. Drawing from episodic memory, the ability to retrieve specific experiences with contexts(Tulving et al., 1972), Zhang et al. (2025a) enable LLMs to learn from episodic experiences for improved planning, Zhang et al. (2025c) integrates episodic memory with semantic memory for LLM personalization, while Chaudhury et al. (2025) introduce episodic attention for processing long contexts.
Reinforcement Learning from Human Feedback (RLHF) A classic behavioral theory, the Operant Conditioning theory posits that behaviors are systematically strengthened or weakened by the consequences (rewards or punishments) that immediately follow them (Skinner, 1957; Thorndike, 1898) principles of reinforcement learning align closely with this psychological framework, particularly in the post-training phase of LLM development, where RLHF explicitly operationalizes Operant Conditioning theory to align model behaviors with human values and preferences. Through repeated feedback, the model gradually adapts to favor outputs that yield higher reward signals—a process akin, in a loose analogy, to Thorndike’s Law of Effect, which describes how behaviors followed by satisfying outcomes tend to recur. While the underlying mechanism is driven by reward optimization algorithms rather than psychological intent, the conceptual resemblance highlights how reinforcement strategies can shape model outputs (Lambert et al., 2023). During RLHF, the model generates responses, and a learned reward function $R(x)$ assigns scores to outputs $x$, guiding subsequent policy updates. For instance, Ouyang et al. (2022) train InstructGPT using Proximal Policy Optimization (Schulman et al., 2017), rewarding responses preferred by humans and penalizing less desirable ones. Foundational frameworks (Christiano et al., 2017; Stiennon et al., 2022; Sutton and Barto, 2018)explicitly translating human judgments into reward signals, operationalizing the insights of Operant Conditioning. More recent work incorporates human cognitive biases (Siththaranjan et al., 2024; Wadi and Fredette, 2025) and personalizes reward functions for individual values (Poddar et al., 2024) and awareness (Liu et al., 2025a). These developments illustrate how Operant Conditioning remains central to aligning LLMs with nuanced human values. While our survey focuses on psychological dimensions, a technical overview of RLHF methods is provided in Appendix LABEL:appendix:rlhf.
@@ 5 Evaluation and Application
Psychology offers tools for both assessing and enhancing model behavior in the evaluation and application stage. We review three key areas of challenges where psychology can inform LLM development: (1) evaluating emergent capabilities such as reasoning, (2) improving task performance in domains involving human cognition, and (3) designing socially aware, multi-agent systems.
@@ 5.1.1 Social Reasoning and Intelligence
Notably, Theory of Mind (ToM) offers a framework for evaluating how individuals understand and attribute mental states – such as beliefs, desires, and intentions – to others. By measuring LLMs’ capacity to reason beliefs, researchers can assess core social intelligence. Recent benchmarks probe different facets of ToM: Chen et al. (2024c); Xu et al. (2024a) provide relatively comprehensive coverage, while others test ToM on higher-order reasoning (Wu et al., 2023), interactions (Kim et al., 2023), white lies (Liu et al., 2025d), cognitive appraisal (Yeo and Jaidka, 2025), and temporally evolving mental states (Xiao et al., 2025). Extending the efforts to spoken dialogues, Soubki et al. (2024) reveal lingering gaps between LLM and human performance. Surveys (Chen et al., 2025b; Ma et al., 2023; Sap et al., 2022; Soubki and Rambow, 2025) consolidate methods and underscore the challenges of robust ToM-based evaluations.
Beyond individual cognition, social influence theories like Conformity Theories (Asch, 2016), capture how group pressure shapes individual judgments. Recent work tests LLM-based agents’ collaboration and bias dynamics under these principles (Choi et al., 2025; Jin et al., 2024b; Wu et al., 2025d; Zhang et al., 2024c; Zhu et al., 2025), bridging individual and group-level cognition.
Emotion is another pillar of social intelligence. Ekman’s Basic Emotion Theory (Ekman, 1992) identifies six universal emotions, often used as labels, while Dimensional Models like the Circumplex Model conceptualize emotions along valence and arousal (Gong et al., 2024; Morrill et al., 2024). LLMs advance on emotion recognition, benefiting dialogue and sentiment tasks (Sabour et al., 2024; Wu et al., 2024c, d; Zhang et al., 2024e).
@@ 5.1.2 Language Proficiency
Recent work adopts psycholinguistic assessments, originally designed for humans, to test LLMs’ language proficiency. These experiments probe a wide range of linguistic domains: morphology (Anh et al., 2024), syntax (Amouyal et al., 2025; An et al., 2025; Hale and Stanojević, 2024; Kandra et al., 2025; Li and Hao, 2025; Liu et al., 2024b), phonology (Goriely and Buttery, 2025; Jang et al., 2025), semantics (Bavaresco and Fernández, 2025; Duan et al., 2025; Hayashi, 2025) and their combinations (Ma et al., 2025a; Miaschi et al., 2024; Zhou et al., 2025a).
Although LLMs exhibit comparable performance to human speakers on many psycholinguistic tasks, the underlying processing mechanism they rely on may seem different from humans (Lee et al., 2024; Pedrotti et al., 2025). Human language acquisition is often characterized by the Poverty of the Stimulus, children acquire complex grammar from relatively little input (Chomsky, 1980), whereas LLMs typically require developmentally implausible amounts of linguistic data to learn morphological rules. On the other hand, some evidence suggests that the learning patterns of LLMs mirror aspects of human language acquisition (Liu et al., 2024b; Zhou et al., 2025c).
Several studies have explored the pragmatic abilities of LLMs, motivated by the close link between language and broader cognitive functions in humans. Grice (1975)’s Theory of conversational implicature posits that utterance interpretation depends on both literal content and surrounding context. Researchers (Bender and Koller, 2020; Gubelmann, 2024) have contrasting perspectives on LLMs with respect to the Harnad (1990)’s Symbol Grounding Problem, i.e. linguistic symbols must be grounded in sensorimotor interactions to be meaningful. Failures of LLMs in pragmatic and semantic tasks (He et al., 2025; Junker et al., 2025; Kibria et al., 2024; Zeng et al., 2025; Ma et al., 2025a), as well as their neuron patterns (He et al., 2025; Wu et al., 2024b), point to limitations beyond pure linguistic knowledge, which potentially parallel human higher-level cognitive processes.